# Minimax Decision Trees via Martingale Approximations

### v1

@@ 2 On partition-based martingale approximations
In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates, thus answering (P-martingale).
2.1 Basic concepts
We say that a sequence of finite partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}^{d}$ is nested if for each $k\geqslant 0$ and $A\in\pi_{k+1}$, there exists (a unique) $B\in\pi_{k}$ such that $A\subseteq B$. Consider nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$ and a real-valued random variable $U$ with a finite second moment. Recall (3):
$M_{k}(\omega):=\mathbb{E}[U\mid U\in A],\quad\text{for }U(\omega)\in A,\leavevmode\nobreak\ A\in\pi_{k},$ (8)
where the MSE $\mathbb{E}[(U-M_{k})^{2}]$ is non-increasing in $k$ by the nested property of $\{\pi_{k}\}_{k\geqslant 0}$ and the total variance formula.
@@ 2.1 Basic concepts
A discrete-time real-valued stochastic process $\{M_{k}\}_{k\geqslant 0}$ is a martingale if for any $0\leqslant\ell<k$, $\mathbb{E}[M_{k}\mid\sigma(M_{0},\dots,M_{\ell})]=M_{\ell}$. We use the abbreviation $\sigma(\pi_{k})$ to denote the $\sigma$-algebra generated by events of the form $\{U\in A\}_{A\in\pi_{k}}$. If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in $\pi_{k}$’s (Doob, 1953). It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)). We call this martingale the $\Pi$-based martingale approximation of the random variable $U$, or in general a partition-based martingale approximation that approximates the random variable in terms of MSE.
@@
Unless $U$ is a constant, there are different kinds of partition, so there are different partition-based martingale approximations depending on the distribution of $U$. Our goal in this section is to identify a few partition-based martingale approximations that efficiently approximate $U$, where the construction algorithm is universal; see (P-martingale). The efficiency criterion is given by the decay of the MSE $\mathbb{E}[(U-M_{k})^{2}]$.
@@
Without loss of generality, we give our construction of a partition of a generic interval $[a,b)\subset\mathbb{R}$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$. [^2] Here we slightly abuse notation that $[a,b)=(-\infty,b)$ if $a=-\infty$. The sequence of partitions $\Pi$ in the response space will be constructed recursively, where $\pi_{0}=\{\mathbb{R}\}$ and for every $k\geqslant 0$, each interval $A\in\pi_{k}$ splits into two intervals, by following the same construction, forming the elements in $\pi_{k+1}$. In the following, we introduce four distinct splitting rules that define partition-based martingale approximations. For simplicity, we assume that $U$ is atomless, so that the endpoints of the intervals do not matter and that no trivial split occurs.
@@
    If the minimizer is not unique, we pick the largest minimizer. The variance splitting rule (corresponding to the VarianceSplit algorithm in Section 3.1) splits $I$ into the two sets $[a,u_{\text{var}})$ and $[u_{\text{var}},b)$.
@@
    Define $u_{\text{Simons}}=\mathbb{E}[U\mid U\in I]$. The Simons splitting rule splits $I$ into the two sets $[a,u_{\text{Simons}})$ and $[u_{\text{Simons}},b)$.
@@
    If the minimizer is not unique, we pick the largest minimizer. The minimax splitting rule (corresponding to the MinimaxSplit algorithm in Section 3.1) splits $I$ into the two sets $[a,u_{\text{minimax}})$ and $[u_{\text{minimax}},b)$.
@@
    The median splitting rule splits $I$ into two sets $[a,u_{\text{median}})$ and $[u_{\text{median}},b)$.
In turn, when applying recursively the variance (resp. Simons, minimax, median) splitting rule starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$. We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp. Simons, minimax, median) martingale (with respect to $U$).
@@
If $U$ is uniformly distributed on a compact interval, all four martingales coincide. For example, if $U\sim\mathrm{Uni}[0,1]$, each of the four martingales from Definition 1 satisfy $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]\asymp 4^{-k}$ (where $\asymp$ means up to universal constants).
@@
In a partition-based martingale, the nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ can be naturally identified as the vertices of a binary tree. A binary tree is the unique infinite tree $(V,E)$ such that all vertices have degree $3$, except for a unique vertex $\emptyset$ called the root, which has degree $2$. Denote by $V_{k}$ the set of all vertices in $V$ with the graph distance from the root equal to $k$. Then, each vertex $v\in V_{k}$ can be identified as a set $A\in\pi_{k}$. With each edge $e=(A_{k},A_{k+1})\in E$ where $A_{k}\in\pi_{k}$ and $A_{k+1}\in\pi_{k+1}$, we may associate a coefficient $p_{e}:=\mathbb{P}(U\in A_{k+1})\in[0,1]$. With each vertex $v=A_{k}$, we may associate a location $\ell_{v}:=\mathbb{E}[U\mid U\in A_{k}]$. It follows that $M_{k}$ is supported on the discrete points $\{\ell_{v}\}_{v\in V_{k}}$ and furthermore,
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}.$ (10)
This binary tree representation (10) of the risk will be frequently used in this paper. In the next two sections, we discuss two types of results: uniform rates for bounded $U$ (Section 2.2) and non-uniform rates (Section 2.3) where the asymptotic constant may depend on the possibly unbounded law of $U$.
2.2 Uniform convergence rates
@@
Let $U$ be a $[0,1]$-valued atomless random variable and $\{M_{k}\}_{k\geqslant 0}$ be one of the four martingale approximations with respect to $U$ given by Definition 1. The following statements hold.
@@
The general case of a bounded $U$ can be derived by a scaling argument, since the constructions in Definition 1 are scale-invariant. Let us sketch the arguments for the minimax martingale, as the same idea becomes crucial when applied to minimax decision trees we develop in Section 3.
Assuming that $U$ is atomless and supported in $[0,1]$, we apply the representation (10). By construction, $\sum p_{(v_{k},v_{k+1})}=1$ and $\sum|\ell_{v_{k}}-\ell_{v_{k+1}}|\leqslant\sup\operatorname{supp}U-\inf\operatorname{supp}U=1$. Moreover, by the minimax property and the law of total variance, it is straightforward to verify (with details in Section 3.2) that the variances decay at least geometrically by half at each split at $u_{\text{minimax}}$, i.e., for any $v\in V_{k}$ with $v^{\prime}\sim w$ for $v^{\prime}\in V_{k-1}$,
$\max_{w\in V_{k+1},\,v\sim w}p_{(v,w)}(\ell_{v}-\ell_{w})^{2}\leqslant\frac{1}{2}p_{(v^{\prime},v)}(\ell_{v^{\prime}}-\ell_{v})^{2}.$
Therefore, by induction,
$\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\leqslant L2^{-k}\mathrm{Var}(U)\leqslant L2^{-k},$
where $L>0$ is a universal constant that may not be the same on each occurrence. By Hölder’s inequality,
$\mathbb{E}[(M_{k}-M_{k+1})^{2}]$ $=\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}$
$\leqslant\bigg{(}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}\bigg{)}^{1/3}$
×$\hskip 28.45274pt\times\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\sum_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}|\ell_{v_{k}}-\ell_{v_{k+1}}|\bigg{)}^{2/3}$
$\leqslant\bigg{(}\max_{\begin{subarray}{c}v_{k}\in V_{k},\,v_{k+1}\in V_{k+1}\\ v_{k}\sim v_{k+1}\end{subarray}}p_{(v_{k},v_{k+1})}(\ell_{v_{k}}-\ell_{v_{k+1}})^{2}\bigg{)}^{2/3}\leqslant L2^{-2k/3}.$
The rest then follows from the martingale property: $\mathbb{E}[(U-M_{k})^{2}]\leqslant\sum_{\ell\geqslant k}\mathbb{E}[(M_{\ell}-M_{\ell+1})^{2}]\leqslant L2^{-2k/3}$.
Finding the optimal rate $r\in(0,1)$ (where $\mathbb{E}[(U-M_{k})^{2}]\leqslant Lr^{k}$ for some universal constant $L$) is also an intriguing question. Observe that one cannot have $r<1/4$ by Example 2 above. In Examples 11 and 12 below, we show that the rate $r$ in Theorem 3 is indeed optimal for the Simons and median martingales.
2.3 Non-uniform convergence rates
The above Theorem 3 requires that the support of $U$ is bounded and the asymptotic constant depends only on the range of $U$, $\sup U-\inf U$. We show in the next result that, without the bounded range assumption, non-uniform convergence rates can still be guaranteed.
@@ 2.3 Non-uniform convergence rates
Let $U$ be an atomless random variable with a finite $(2+\varepsilon)$-th moment for some $\varepsilon>0$, and $\{M_{k}\}_{k\geqslant 0}$ be one of the four partition-based martingale approximations given by Definition 1. Then there exist constants $r\in(0,1)$ and $C>0$, both possibly depending on the law of $U$, such that
@@
If $U$ is bounded, Theorem 4 is a special case of Theorem 3. The case of the Simons martingale has been established by Zhang et al. (2024), and our proof of Theorem 4 will follow a similar route. Since the asymptotic constant is allowed to depend on the law of $U$, the optimal non-uniform rate $r$ might be smaller than the uniform ones (in Theorem 3).
Recall from Section 2.2 that the rates obtained in Theorem 3 are asymptotically optimal for the Simons and median martingales (see Examples 11 and 12). For the variance and minimax martingales, we show in the next result that the rate $r=1/4$ is optimal under certain regularity conditions on the law of $U$.

### v2

@@ Appendix A On partition-based martingale approximations
In this section, we develop the framework of partition-based martingale approximations and show that a number of constructions feature exponential convergence rates. These problems are of their own interest (see the end of Appendix A.1) and have implications in the regression tree setting. The main takeaway is that:
  • The techniques developed in this work for analyzing convergence rates of the MinimaxSplit algorithm extend to many other settings, such as convergence rates for martingale approximations;
  • Precise (asymptotic) convergence rates of martingale approximations depend on the initial law, but a uniform exponential rate is always guaranteed under mild conditions;
  • Applying VarianceSplit to samples $\{(\mathbf{X}_{i},Y_{i})\}_{1\leqslant i\leqslant n}$ satisfying monotonicity in all $d$ dimensions (meaning that for every $j\in[d]$ and distinct $i,i^{\prime}\in[n]$, $((\mathbf{X}_{i})_{j}-(\mathbf{X}_{i^{\prime}})_{j})(Y_{i}-Y_{i^{\prime}})=\sigma_{j}$ where $\sigma_{j}\in\{\pm 1\}$ depends only on $j$) does not incur ECP.
A.1 Basic concepts
@@
Consider nested partitions $\{\pi_{k}\}_{k\geqslant 0}$ of $\mathbb{R}$ and a real-valued random variable $U$ with a finite second moment. We use the abbreviation $\sigma(\pi_{k})$ to denote the $\sigma$-algebra generated by events of the form $\{U\in A\}_{A\in\pi_{k}}$. Define $M_{k}:=\mathbb{E}[U\mid\sigma(\pi_{k})]$. If $\Pi:=\{\pi_{k}\}_{k\geqslant 0}$ is a nested sequence of partitions, $\{\sigma(\pi_{k})\}_{k\geqslant 0}$ is a filtration generated by indicator functions of sets in $\pi_{k}$’s (Doob, 1953). It follows from the tower property of conditional expectations that the sequence $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ is a martingale (in fact, a Doob martingale (Doob, 1940, 1953)). We call this martingale the $\Pi$-based martingale approximation of the random variable $U$, or in general a partition-based martingale approximation that approximates the random variable $U$ in $L^{2}$.
@@
Unless $U$ is a constant, there are different partition-based martingale approximations depending on the distribution of $U$. Our goal in this section is to identify a few partition-based martingale approximations that efficiently approximate $U$, where the construction algorithm is universal. The efficiency criterion is given by the decay of the MSE $\mathbb{E}[(U-M_{k})^{2}]$. The MSE is non-increasing in $k$ by the nested property of $\{\pi_{k}\}_{k\geqslant 0}$ and the total variance formula.
@@
Without loss of generality, we give our construction of a partition of a generic interval $a,b)\subset\mathbb{R}$, where $a,b\in\mathbb{R}\cup\{\pm\infty\}$. [^8] Here we slightly abuse notation that $a,b)=(-\infty,b)$ if $a=$-$\infty$. The sequence of partitions $\Pi$ will be constructed recursively, where $\pi_{0}=\{\mathbb{R}\}$ and for every $k\geqslant 0$, each interval $A\in\pi_{k}$ splits into two intervals, by following the same construction, forming the elements in $\pi_{k+1}$. In the following, we introduce four distinct splitting rules that define partition-based martingale approximations. For simplicity, we assume that $U$ is atomless, so that the endpoints of the intervals do not matter and that no trivial split occurs. The general case with atoms can be analyzed in a similar way but with more technicality.
@@
    If the minimizer is not unique, we pick the largest minimizer. The variance splitting rule (corresponding to the VarianceSplit algorithm in Section 2.1) splits $I$ into the two sets $a,u_{\text{var}})$ and $u_{\text{var}},b)$.
@@
    Define $u_{\text{Simons}}=\mathbb{E}[U\mid U\in I]$. The Simons splitting rule splits $I$ into the two sets $a,u_{\text{Simons}})$ and $u_{\text{Simons}},b)$.
@@
    If the minimizer is not unique, we pick the largest minimizer. The minimax splitting rule (corresponding to the MinimaxSplit algorithm in Section 2.1) splits $I$ into the two sets $a,u_{\text{minimax}})$ and $u_{\text{minimax}},b)$.
@@
    The median splitting rule splits $I$ into two sets $a,u_{\text{median}})$ and $u_{\text{median}},b)$.
In turn, when applying recursively each splitting rule starting from $\pi_{0}=\{\mathbb{R}\}$, we obtain a nested sequence of partitions $\Pi=\{\pi_{k}\}_{k\geqslant 0}$ depending on the law of $U$. We call the resulting $\Pi$-based martingale approximation $\{M_{k}\}_{k\geqslant 0}=\{\mathbb{E}[U\mid\sigma(\pi_{k})]\}_{k\geqslant 0}$ the variance (resp. Simons, minimax, median) martingale (with respect to $U$). See Figure 10 for examples of the four martingale approximations and their convergence rates.
Figure 10: Plots of the log-MSE, $\log_{2}(\mathbb{E}[(U-M_{k})^{2}])$ versus the approximation depth $k$ for four different methods, where the density of the law $U$ is given: (a) by $f(u)\propto{\mathbbm{1}}_{[0,0.9]}(u)+(1+10^{4}(u-0.9)){\mathbbm{1}}_{[0.9,1]}(u)$ and (b) by $f(u)\propto u^{10}{\mathbbm{1}}_{[0,1]}(u)$. The right panel of (b) shows the ratios $\mathbb{E}[(U-M_{k+1})^{2}]/\mathbb{E}[(U-M_{k})^{2}]$ for the four methods as a function of $k$.
@@
If $U$ is uniformly distributed on a compact interval, all four martingales coincide. For example, if $U\buildrel\mathrm{law}\over{\sim}\mathrm{Unif}(0,1)$, each of the four martingales from Definition 14 satisfy $M_{k}=\mathbb{E}[U\mid\sigma(\pi_{k})]$, where $\pi_{k}:=\{[j/2^{k},(j+1)/2^{k}):0\leqslant j<2^{k}\}$ and it follows that $\mathbb{E}[(U-M_{k})^{2}]=4^{-k}/12$.
@@
Motivated by a martingale embedding problem, Simons (1970) first introduced the Simons martingale and established the a.s. convergence $M_{k}\to U$ (and hence also in $L^{2}$), but did not analyze the convergence rate. A recent motivation for studying the convergence rate of the Simons martingale arises from the construction of powerful e-values in hypothesis testing (Ramdas and Wang, 2025). Zhang et al. (2024) (Lemma 5.6) proved that if $U\in L^{2+\delta}$ for some $\delta>0$ and $\{M_{k}\}$ is the Simons martingale, then there exist $C>0$ and $r\in(0,1)$ such that $\mathbb{E}[(U-M_{k})^{2}]\leqslant Cr^{k}$, and $r<0.827$ is feasible if $U$ is bounded. Our Theorem 16(ii) improves the rate to $r=1/2$ and hence provides a tighter theoretical bound. We also show by example that the rate $r=1/2$ is optimal.
The terminology martingale approximation has been used extensively in the probability literature with different meanings. In Rüschendorf (1985), it refers to the best approximation of a random vector by a (single) martingale based on ideas from optimal transport. In Borovskikh and Korolyuk (1997) and Hall and Heyde (2014), it refers to techniques from martingale theory (such as inequalities and CLT rates) with various applications in statistics. Similar techniques are also used in the study of stationary ergodic sequences (Wu and Woodroofe, 2004; Zhao and Woodroofe, 2008) and Markovian walks (Grama et al., 2018). Note that in our setting, $Y$ is a sum of martingale differences, but whose variance is bounded, and hence one cannot apply the martingale CLT or its convergence rates.
In the next two sections, we discuss two types of results: uniform rates for bounded $U$ (Appendix A.2) and non-uniform rates (Appendix A.3) where the asymptotic constant may depend on the possibly unbounded law of $U$.
A.2 Uniform convergence rates
@@
Let $U$ be a $[0,1]$-valued atomless random variable. The following statements hold.
@@
The proof of Theorem 16 is lengthy and technical, which we defer to Appendix A.4.
The general case of a bounded $U$ can be derived by a scaling argument, since the constructions in Definition 14 are scale-invariant. Finding the optimal rate $r\in(0,1)$ (where $\mathbb{E}[(U-M_{k})^{2}]\leqslant Lr^{k}$ for some universal constant $L$) is also an intriguing question. Observe that one cannot have $r<1/4$ by Example 15 above. In Examples 17 and 18 below, we show that the rate $r$ in Theorem 16 is indeed optimal for the Simons and median martingales.
Example 17 (Rate $r=1/2$ is optimal for the Simons martingale).
$\mathbb{P}\bigg{(}U=\sum_{j=1}^{J}\Big{(}\frac{1-s}{s}\Big{)}^{j}\bigg{)}=s^{J+1}(1-s).$
It follows that $U$ is a bounded random variable. Next, we compute that
$\mathbb{E}\bigg{[}U\mid U\geqslant\sum_{j=1}^{J}\Big{(}\frac{1-s}{s}\Big{)}^{j}\bigg{]}$ $=\bigg{(}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\bigg{)}^{-1}\sum_{\ell=J}^{\infty}s^{\ell+1}(1-s)\sum_{j=1}^{\ell}\Big{(}\frac{1-s}{s}\Big{)}^{j}$
$=s^{-(J+1)}\sum_{\ell=J}^{\infty}s^{\ell+1}(2s-1)\bigg{(}1+\Big{(}\frac{1-s}{s}\Big{)}^{\ell}\bigg{)}$
$=\frac{s^{-(J+1)}(2s-1)s^{J+1}}{1-s}+\frac{s^{-(J+1)}(2s-1)s}{1-(1-s)}$
$=\sum_{j=1}^{J+1}\Big{(}\frac{1-s}{s}\Big{)}^{j}.$
$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=(1-s)^{2k+1}s^{-(k+1)}\asymp\Big{(}\frac{(1-s)^{2}}{s}\Big{)}^{k}.$
Example 18 (Rate $r=1/2$ is optimal for the median martingale).
Let $s\in(0,1)$ and a bounded random variable $U$ satisfy
Next, we compute that
$=2^{k}\sum_{\ell=k+1}^{\infty}2^{-\ell}\Big{(}\frac{1}{1-s}-\frac{s^{\ell-1}(2-s)}{1-s}\Big{)}$
$=\frac{1}{1-s}-\frac{2-s}{s(1-s)}\frac{s^{k+1}/2}{1-s/2}$
$=\sum_{j=1}^{k}s^{j-1}.$
Note also that
$\mathbb{E}[(M_{k+1}-M_{k})^{2}]=2^{-k}s^{2k}.$
A.3 Non-uniform convergence rates
The above Theorem 16 requires that the support of $U$ is bounded and the asymptotic constant depends only on $\sup\operatorname{supp}U-\inf\operatorname{supp}U$. We show in the next result that, without the bounded range assumption, non-uniform convergence rates can still be guaranteed.
@@ A.3 Non-uniform convergence rates
Let $U$ be an atomless random variable with a finite $(2+\varepsilon)$-th moment for some $\varepsilon>0$, and $\{M_{k}\}_{k\geqslant 0}$ be one of the four partition-based martingale approximations given by Definition 14. Then there exists a constant $r\in(0,1)$ depending only on $\varepsilon$ and another constant $C>0$ depending on the law of $U$, such that
@@
Proof.
Let $\operatorname{supp}M_{1}=\{x,y\}$ where $x<y$. Take $M>\sup\{x,y\}$. Let $\tau_{M}^{+}$ be the first hitting time to $\{x:x\geqslant M\}$, $\tau_{M}^{-}$ be the first hitting time to $\{x:x\leqslant-M\}$, and $\tau_{M}=\min\{\tau_{M}^{+},\tau_{M}^{-}\}$. It is straightforward to prove (see the proof of Lemma 5.6 of Zhang et al. (2024)) that for some $r\in(0,1)$ and $C>0$,
Similarly, the same analysis holds for $\tau_{M}^{-}$. Hence, we conclude that $\mathbb{P}(\tau_{M}<\infty)=O(M^{-1}).$ The rest follows line by line as in Zhang et al. (2024), which we sketch for completeness. By Hölder’s inequality, $\mathbb{E}[{\mathbbm{1}}_{\{\tau_{M}<\infty\}}U^{2}]=O(M^{-\delta^{\prime}})$ for some $\delta^{\prime}>0$ depending only on $\varepsilon$. With the choices $M=r^{-k/(2+\delta^{\prime})}$ and $q=r^{\delta^{\prime}/(2+\delta^{\prime})}\in(0,1)$, we have that for some $C>0$,
If $U$ is bounded, Theorem 19 is a special case of Theorem 16. The case of the Simons martingale has been established by Zhang et al. (2024), and our proof of Theorem 19 follows a similar route. Since the asymptotic constant is allowed to depend on the law of $U$, the optimal non-uniform rate $r$ might be smaller than the uniform ones (in Theorem 16).
Recall from Appendix A.2 that the rates obtained in Theorem 16 are asymptotically optimal for the Simons and median martingales (see Examples 17 and 18). For the variance and minimax martingales, we show in the next result that the rate $r=1/4$ is optimal under certain regularity conditions on the law of $U$.
@@
Proof.
To see this, without loss of generality we start from $U$ supported on $[0,1]$ (meaning that $\mathbb{P}(U\in[0,1])=1$) whose density takes values in $1,1+\varepsilon]$ (we may normalize afterwards to a probability measure, which we omit). Then for any interval $a,b]\subseteq[0,1]$,
Therefore,
$\leqslant\frac{2(1+\varepsilon)^{4}}{3(2+\varepsilon)^{3}}(b-a)^{3}.$
Similarly,
It follows that the split point $x_{*}\in[0,1]$ must satisfy
Hence, as $\varepsilon\to 0$, the ratio becomes
as desired. ∎